OpenStudy (anonymous):

Simplify the following expression. \[\sqrt{-18}\sqrt{-72}\]

OpenStudy (lgbasallote):

\[\sqrt{(-18)(-72)}\] can you use a calculator?

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

I DONT HAVE ONE

OpenStudy (anonymous):

HELP PLEASE

OpenStudy (chris):

I think the question was more - are you allowed to on this problem.

OpenStudy (lgbasallote):

okay..i'll do this the manual way...

OpenStudy (lgbasallote):

yuppp that's what i meant thanks @chris :D

OpenStudy (chris):

np =). and I think we're assuming not allowed heh.

OpenStudy (zarkon):

\[\sqrt{-18}\sqrt{-72}\neq \sqrt{(-18)(-72)}\]

OpenStudy (lgbasallote):

oh man..made a mistake in one :O

OpenStudy (anonymous):

WHAT WAS THE MISTAKE

OpenStudy (lgbasallote):

isn;t it @Zarkon ? isn't \[\sqrt{a}\sqrt{b} = \sqrt{(a)(b)}\]?

OpenStudy (zarkon):

if \(a,b\ge 0\)

OpenStudy (lgbasallote):

oh..hmm..yeah..this is imaginary now that i think about it..

OpenStudy (zarkon):

\[\sqrt{-18}\sqrt{-72}\] \[=i\sqrt{18}\times i\sqrt{72}\] \[=i^2\sqrt{18}\times \sqrt{72}\] \[=-\sqrt{18\times 72}\] \[=-\sqrt{18\times 2\times 36}\] \[=-\sqrt{36\times 36}=-36\]

OpenStudy (lgbasallote):

hmm yeah..went pretty careless there..sorry :)

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

THANKS EVERYONE

OpenStudy (lgbasallote):

\[\sqrt{-a} = i \sqrt{a}\] he just factored an i..then when i^2..it became -1...my other mistake was the 2 i pulled out..i wrote 4 so my answered doubled..do you get it now?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):